The sum of the three angles in any triangle is 180 degrees. The smallest angle in a given triangle is one-half the second angle and 52 degrees smaller than the largest angle. How many degrees does each angle contain?
Please, I just need the formula set up so I can try to understand how to get it.

Answer by dolly(163)

Let the smallest angle be x
then the second angle = 2x
The largest angle = x + 52
Sum of all the angles = 180
==> x + 2x + x + 52 = 180
==> 4x + 52 = 180
==> 4x = 180 - 52
==> 4x = 128
==> x = 128/4
= 32
Thus the angles of the triangle are 32, 64, 84 degrees